A Complete Bibliography of Applied Statistics: 2020–2029

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 October 2023 Version 1.06

Title word cross-reference

χ^{2} [36]. G [89, 58]. K [156, 142]. M [50].

-computation [58, 89]. -means [156]. -quantile [50]. -sample [142]. -test [36].

1860s [199]. 19 [178, 198].

abilities [93, 78]. abundance [129]. Accelerating [112]. accents [176]. accidents [74]. accommodating [73]. Accounting [129, 19, 224]. across [147]. activity [98]. actors [129]. adaptive [26, 131]. Adding [61]. additive [48]. Adjusting [106]. administrative [102]. adolescent [152]. adoptive [110]. advanced [229, 212]. adverse [152]. Africa [165]. after [189]. again [222]. age [223, 198]. agent [128]. agents [226]. aggregation [79]. AIDS [171]. air [174, 146, 169, 35, 209]. airport [107]. alignment [16]. allowing [61]. Alzheimer [73]. analyses [54]. Analysing [184, 9]. Analysis [2, 14, 183, 22, 45, 60, 62, 135, 103, 176, 91, 83, 88, 54, 185, 199, 203, 201, 133, 116, 136, 122, 58, 57, 80, 166, 218, 126, 63, 82, 142]. animal [204]. animals [182]. annual [115]. anticoagulant [26]. application [76, 137, 11, 15, 32, 60, 55, 104, 71, 66, 9, 18, 179, 159, 194, 123, 118, 8, 116, 50, 122, 70, 41, 166, 35, 154, 215, 209, 79, 63, 82, 97, 145]. applications [187, 142, 171]. approach [95, 32, 208, 174, 101, 197, 210, 189, 195, 58, 42, 41, 126, 78]. approaches [145]. approximate [222]. approximately [216]. area [115]. areal [101, 115]. arms [114]. arrival [202]. artificial [185]. Assessing [119, 21, 173, 98, 192, 215]. Assessment [50]. assignment [34]. association [95, 90, 207]. attribute [207]. augmentation [15, 148]. augmented [24]. authors [67]. autism [88]. automatic [104]. autonomous [185]. autopsy [96]. autoregressive [205, 164]. average [130]. averaging [127]. avian [133].

bagging [72]. bank [97]. based [229, 40, 102, 119, 208, 84, 212, 111, 48, 123, 158, 190, 97]. basin [39]. basket [227]. basketball [65]. bath [176]. batsmen [93]. Bayesian [137, 229, 183, 113, 34, 220, 112, 64, 16, 62, 152, 101, 19, 128, 71, 225, 134, 178, 89, 17, 169, 207, 212, 9, 92, 111, 24, 125, 194, 96, 227, 23, 199, 8, 77, 201, 133, 153, 116, 136, 7, 126, 78, 165, 190, 209, 211, 82]. behavioural [182]. behaviours [152]. Beijing [169]. being [201]. between [196, 105, 59]. Beyond [135]. bias [54]. bike [80]. binary [28, 48]. binomial [148]. biomarkers [5, 192]. biomedical [10]. birth [35]. bivariate [90]. bleeding [87]. block [99, 145]. blood [190, 63]. borrowing [130, 227]. boundaries [216, 147]. Bradley [153]. brain [170, 88, 132]. breaking [147]. breast [229, 212, 82]. British [176]. building [86, 5]. Burglary [53]. business [85].

calendar [223]. calibrated [63]. calibration [223]. cancer [229, 187, 128, 212, 82]. capture [21]. card [97]. cardio [181]. cardio-mechanic [181]. Cardiovascular [9]. Carlo [211]. case [72, 24, 115, 107]. case-study [24]. categorical [104, 70]. Causal [39, 89, 58]. cause [96]. cell [110]. Censored [114, 186, 158, 58]. censoring [72, 192]. chain [122]. change [211]. changepoint [91, 98]. changing [144]. Characteristic [141]. characteristics [44]. checks [83]. chemotherapeutic [128]. child [70]. children [159, 63]. choice [222, 122]. Christofides [207]. churn [218]. Circular [66]. city [80]. class [60, 215]. Classification [190, 60]. clean [42]. clinical [110, 128, 226, 171]. Cluster [57, 59, 143]. cluster-randomized [143]. clustered [20]. Clustering [104, 97, 32, 209]. clusters [172]. coarseness [4]. coastal [91]. coefficient [137, 71]. coefficients [119]. coffee [222]. Cohort [76, 89]. cointegration [24]. Combining [226, 128, 181, 115]. comparison [144, 149]. comparisons [155]. compensation [208]. completion [220]. complex [120, 47]. components [208]. composite [183]. compositional [166]. compression [195]. computation [89, 58]. computationally [113]. computer [179]. concordance [119]. concurrent [29]. conditional [215]. confidence [30]. confidentiality [61]. congestion [180]. connection [170]. consistent [61]. constraints [18]. constructing [154]. contact [143]. contagion [172]. Contents [68, 228, 138]. continuing [189]. continuous [52, 135, 226, 122]. continuous-time [122]. Contour [216]. contraceptives [15]. contrast [190]. contrast-enhanced [190]. control [26]. cooking [42]. copula [135, 169, 9, 48]. copula-based [48]. copulas [44]. corporate [172]. Correcting [77]. corrections [43]. correlation [119]. correspondence [203]. Corrigendum [137]. costs [58]. count [168, 55, 116]. counterfactual [123]. counts [178]. coupled [168]. covariance [145]. covariances [146]. covariate [209]. covariate-dependent [209]. covariates [184, 73, 164, 10, 41]. COVID [178, 198]. COVID-19 [178, 198]. Cox [150, 20]. credibility [167]. credit [8]. cricket [93]. criterion [111]. criterion-based [111]. crowdsourced [77]. crush [24]. cumulative [56]. curve [23, 155]. curves [210]. customer [218]. cycling [184]. cytotoxic [226].

Daily [146, 98]. dairy [32]. Danube [39]. data [76, 52, 14, 15, 130, 186, 32, 55, 101, 120, 19, 28, 148, 184, 21, 61, 103, 90, 117, 176, 181, 127, 20, 88, 105, 156, 159, 185, 129, 96, 161, 143, 158, 193, 10, 77, 116, 57, 218, 165, 209, 200, 97, 206]. death [89, 96]. deaths [198]. decentralized [47]. decision [15]. deconvolution [183]. deep [46, 165]. deep-sequence [165]. default [100, 6]. defaults [172]. delivery [126]. dependence [215]. dependent [209, 192]. dependently [158]. deprivation [153]. depth [186]. Derivation [224]. derivatives [23]. design [86, 128, 207, 31]. designs [84, 99, 171, 222]. detecting [147]. detection [91, 195, 98]. determinations [223]. determining [78]. develop [20]. Development [220]. diachronic [211]. diagnostic [83]. dietary [224]. differences [106, 196, 194, 23]. different [4]. Difficulties [159]. diffusion [10]. dimensional [113, 103, 179, 156, 118]. dimensionally [160]. direction [66]. Dirichlet [57]. Dirichlet-multinomial [57]. disability [8]. discharges [39]. discovering [196]. discovery [113, 156]. discrete [149, 147]. discrimination [192]. disease [168, 73, 156, 193]. disequilibrium [118]. disposal [147]. dissolution [155]. distributed [35]. distribution [217]. distributional [135]. DNA [120, 19]. doctor [48]. domain [221]. domains [85]. dose [226]. driven [205, 90]. dropout [89]. drug [27]. during [199]. dyadic [38]. Dynamic [130, 73, 193, 107, 174, 108, 56, 3, 70, 190]. dynamically [144]. dynamics [205, 225, 83].

early [189]. echo [160]. ecological [77].

economic [125]. effect

[55, 18, 48, 143, 189, 126, 63]. effects [52, 174, 225, 61, 221, 197, 10, 136, 50, 70]. efficacy [41]. efficiency [214]. Efficient [15, 219, 113]. electronic [177]. employee [208]. employment [102]. emulation [179]. enclosing [216]. endogenous [48]. energy [42]. Engel [210]. engines [44]. England [198]. enhanced [190]. ensemble [73, 79]. Environmental [210, 19]. epidemics [157]. epidemiology [187]. equation [174]. equidistant [173]. Erratum [162]. error [61]. errors [19, 77]. eruptions [7]. estimate [174, 143]. estimates [112]. Estimating [6, 17, 140, 48, 126, 220, 189, 93, 42]. Estimation [102, 18, 145, 219, 112, 107, 61, 85]. estimator [218]. ETAS [202]. ethnic [194]. European [38, 13]. Evaluation [167]. event [9, 213]. events [219, 140, 9, 185, 87]. evidence [38, 171]. ex [181]. example [180]. excess [187]. Exchange [11]. exchanges [70]. experiment [222]. experiments [134, 99, 122]. exploratory [117]. Exploring [176, 114, 204]. exposure [61, 59, 63]. extent [13]. extreme [40, 91, 39, 13, 157]. extremes [180].

facility [47]. Factor [24, 186, 96].

Factor-augmented [24]. factors [55]. false [19]. famine [199]. Fault [47]. features [173, 182]. feet [93]. field [115]. fields [191]. final [100]. financial [11]. Finding [93, 64]. findings [36]. fingerprints [141]. Finite [82]. Finnish [199]. fitting [173]. Flexible [221, 22, 187]. flow [107, 165, 190]. flows [80]. flu [180]. flume [122]. football [78]. Forecasting [160, 66]. forecasts [79]. forensic [60, 46, 114]. forests [49, 221, 66]. fracturing [147]. framework [22, 187]. free [95, 107]. frequency [160, 83, 125]. frequentist [112]. fuel [44]. fuels [42]. full [8]. function [80]. function-on-function [80]. Functional [73, 103, 137, 32, 29, 71, 184, 176, 177, 195, 3, 209]. functions [56]. Fused [132]. fusion [108]. Future [86]. futures [11, 121].

games [64]. gamma [148]. gas [11]. Gaussian [52, 186, 150, 155, 191, 93, 200]. gene [45, 82]. general [15]. generalised [207]. Generalized [33, 31, 177, 8]. Generalizing [171]. genomic [118]. geochemical [186]. geolocation [46]. geological [173]. geostatistical [115]. Gibbs [217]. Global [46, 42]. Goal [220]. goats [32]. graphical [4, 132, 200]. graphs [194]. Greenland [17]. grocery [167]. group [34, 31]. growth [23]. guided [156].

haemophilia [87]. hand [63]. harmonic [36]. Hawkes [74, 204]. hazard [187]. hazards [127, 56, 41]. Health [9, 174, 152, 55, 101, 177, 126]. height [120, 18]. heterogeneity [21, 92, 224]. Heterogeneous [200, 53, 206]. hidden [159]. hierarchical [55, 17, 9, 31, 96, 23, 42, 190]. hierarchy [204]. high [113, 160, 179, 156, 118]. high-dimensional [113, 179, 156, 118]. high-frequency [160]. Highly [177, 133]. historical [130]. history [86]. HIV [165]. home [98]. homestead [123]. homogeneous [149]. hospital [213, 180]. household [42]. human [174]. hurdle [55]. hybrid [41]. hydraulic [147]. hydrocracking [134]. hypothesis [88].

I-II [226]. I/II [128]. identifying [55]. ignorable [102]. II [128, 226]. imaging [190]. immunotherapeutic [128]. impact [100]. impacting [55]. imperfect [118]. implementation [8]. improve [143, 214].

Improving [181]. imputation [28, 178]. incidence [56, 105]. incomplete [7]. independently [158]. Index [67]. indicator [220]. indices [108]. indirect [136]. individualized [154]. induced [192]. InfCareHIV [72]. infection [165]. infectious [168]. infer [96]. Inference [40, 120, 10, 49, 150, 89, 181, 18, 129, 211]. Inferring [90, 165]. inflated [20, 116]. influenza [225, 143, 133, 157]. informal [102]. Information [37, 51, 69, 81, 94, 109, 124, 139, 151, 163, 175, 188, 105, 227, 1, 12, 25]. informative [59]. injector [44]. innovation [196]. insights [53]. inspired [86]. insurance [48]. integer [164]. integer-valued [164]. integrating [156]. integration [206]. intelligence [185]. interaction [217, 129]. interactions [65, 204]. interbank [38]. International [50, 107, 8]. interplay [59]. interpretability [167]. intervention [45]. interventions [126]. inverse [72]. Investigating [207]. irregular [177]. isolation [47]. Issue [37, 51, 69, 81, 94, 109, 124, 139, 151, 163, 175, 188, 1, 12, 25]. item [196].

jet [44]. joint [217, 30, 9, 4, 193].

kernel [5, 147, 82]. kinetics [32].

Kolmogorov [97]. Korea [196]. kriging [134].

labelling [104]. lag [178, 35]. Landmark [56]. large [145]. lasso [4, 132]. latent [38, 148, 43, 204, 215]. lead [63]. learning [106, 134]. least [206]. length [126]. lengths [213]. level [86, 225]. level-set [86]. levels [226, 91, 63]. Leveraging [214]. lifetime [140, 127]. Lifting [161]. Likelihood [107]. Likelihood-free [107]. Linear [52, 29, 108, 169, 20, 177, 118, 33]. lines [173]. linkage [118, 215]. location [61]. loci [113]. log [104, 150]. logistic [84, 184]. London [53]. Long [91]. Long-term [91]. Longitudinal [38, 3, 76, 159, 193]. loss [14, 100]. low [6, 35]. low-default [6].

M [76]. M-quantile [76]. machine [106, 82]. machines [5]. magnetic [190]. main [202]. main-shock [202]. making [15]. malaria [105]. manifolds [33]. MAP [197]. Mapping [105, 137, 71]. marginal [219]. marine [104]. mark [41]. mark-specific [41]. marked [7]. market [11, 38, 197]. markets [199]. Markov [168, 159, 191, 65, 122]. marks [41]. massive [218]. matched [63]. matching [86]. maternal [224]. mathematics [50]. matrices [145]. matters [100]. mean [219, 36, 140]. means [156]. measurement [52, 61]. measurements [119]. measures [83, 70]. Measuring [211]. mechanic [181]. mechanism [102, 39]. mediation [136]. medical [58]. membership [101]. Mercantile [11]. messages [104]. meta [54]. meta-analyses [54]. metabolic [194]. method [144]. methods [106, 207, 211]. microbiome [119, 116, 57]. migrations [49]. Milan [80]. milking [32]. Millennium [76]. minutiae [141]. misaligned [101]. misclassification [59, 77]. missing [28, 129, 41]. mitochondrial [142]. Mixed [125, 52, 55, 135, 225, 221, 18, 159, 10]. mixed-effects [225]. Mixed-frequency [125]. mixtures [120, 57, 82]. MNAR [89]. mobility [80]. model [137, 183, 113, 95, 186, 220, 64, 208, 217, 152, 55, 71, 225, 74, 169, 9, 92, 127, 179, 158, 164, 115, 8, 153, 93, 202, 42, 41, 35, 147, 200]. model-based [208]. model-free [95]. modeling [155]. Modelling [172, 11, 44, 19, 205, 180, 13, 80, 22, 187, 101, 134, 196, 17, 114, 213, 182, 194, 227, 23, 193, 65, 153, 87, 7, 176]. models [52, 172, 167, 15, 216, 112, 168, 135, 148, 84, 108, 18, 56, 24, 48, 159, 96, 43, 191, 118, 10, 33, 122, 70, 57, 204, 165, 215, 190, 211]. Monte [211]. morbidity [146]. mortality [146]. mortality/morbidity [146]. motorway [74]. movement [182]. MRI [166]. multi [183, 186, 225, 145]. multi-depth [186]. multi-level [225]. multi-omic [145]. multi-region [183]. Multifidelity [179]. Multilevel [196, 28]. multinomial [57]. Multiple [149, 28, 34, 152, 101, 223, 73, 182, 194, 209]. multiple-assignment [34]. Multiscale [88]. Multisensor [108]. multistate [21]. multivariate [76, 15, 150, 148, 146, 117, 159, 191, 116, 70, 42, 209]. myocardial [190].

n [34]. natural [11, 136]. nearest [186]. nearest-neighbour [186]. necessary [141]. negative [19, 200]. neighbour [186]. Nested [58, 171]. Netherlands [133]. Network [204, 112, 152, 196, 88, 210, 39, 129, 214]. network-valued [88]. networks [11, 38, 107, 46, 160, 99, 88, 161, 132]. neural [46, 210]. no [6]. no-default [6]. Non [223, 191, 209, 52, 102, 74, 169, 171, 200]. non-Gaussian [52, 200]. non-ignorable [102]. non-linear [169]. non-negative [200]. non-nested [171]. Non-parametric [223, 209, 74]. Non-separable [191]. nonlinear [83]. Nonparametric [142, 92, 194, 133, 116]. novel [32]. Nowcasting [198, 178, 125]. nuclear [140]. null [88]. number [134].

observational [63]. observations

[104, 115]. observed [140]. observing [189]. Ohio [147]. omic [145]. omics [206]. One [60, 218]. One-class [60]. one-step [218]. Optimal [99, 222, 5, 31]. order [150]. ordered [135]. ordered-continuous [135]. ordinal [184]. origin [173]. outbreak [133]. Outcome [156]. Outcome-guided [156]. outcomes [135, 189]. output [179]. overall [30, 143]. paired [144, 117]. pairwise [217]. panel [15, 195, 70]. parallel [173]. parameter [107]. parameters [144]. parametric [22, 223, 150, 89, 74, 213, 164, 165, 209]. parent [70]. partial [8, 206]. partially [140, 33]. passenger [107]. pathogenic [133]. pathway [82]. patient [126]. patients [88]. pattern [92]. patterns [2, 224, 98]. peak [120]. penalization [136]. penalized [27]. performance [65, 192]. period [98]. personalized [229, 212]. perspective [77]. phase [128, 84, 226, 110]. phenotypes [170]. phenotypic [156]. physical [216]. physics [140]. placement [147]. player [78]. PM2.5 [200]. PO2PLS [206]. point [217, 2, 53, 115, 7]. Poisson [112, 165]. policies [197]. policy [123]. pollutants [209]. polluting [42]. pollution [174, 169, 35]. Pólya [148]. Pólya-gamma [148]. polygons [216]. pooled [214]. population [106]. populations [30, 171]. portfolios [6]. positive [19, 178]. Posterior [167]. poverty [135]. power [121, 160]. practitioner [15]. predict [208]. Predicting [170, 64]. Prediction [118, 11, 73, 221, 127, 56, 123, 157, 27]. predictions [115]. predictive [125, 192]. presence [58, 192]. pressure [181]. prevalence [105]. Pricing [121]. primary [74]. principal [189]. probabilistic [66, 79, 206]. probability [6, 72]. probit [15]. procedures [149]. process [186, 217, 134, 74, 155, 93, 202, 204]. processes [150, 53, 7]. production [17]. professional [55]. profit [14]. profit-and-loss [14]. program [143]. Programme [50]. progression [73]. proofing [86]. propensity [21]. proportional [56, 41]. proposal [201]. protect [61]. protein [16]. public [123]. publication [54]. pup [17].

quality [146, 145]. Quantifying [75]. Quantile [83, 159, 76, 50].

Quantile-frequency [83]. quantitative [113]. quest [64]. Questionnaire [159]. queueing [107]. quotient [23].

radial [183]. radiocarbon [223]. Random [70, 49, 28, 221, 207, 191, 50]. random-effects [50]. randomised [207, 227]. randomized [34, 143]. rank [117, 20]. Ranking [154, 30]. rankings [205]. rates [100, 220]. Real [157]. Real-time [157]. recapture [21]. record [215]. records [102, 177, 7]. Recurrent [87, 219, 9, 185]. reduce [172]. reduced [160, 117]. reduced-rank [117]. Reducing [134]. Regime [172]. Regime-switching [172]. region [183, 30, 198]. regional [224]. Register [72]. regression [76, 29, 84, 184, 117, 66, 114, 127, 18, 20, 48, 96, 177, 158, 50, 3, 57, 80, 166, 27]. regressions [113, 82]. regular [196]. Regularized [166, 32]. related [223]. relationships [38, 120]. Reliability [185]. remotely [108]. renewal [202]. repeated [52, 70]. replicated [2]. reporting [178]. reports [2]. reproducibility [119]. required [134]. rerandomization [131]. reservoir [183]. resolution [100]. resonance [190]. respondent [90]. respondent-driven [90]. Response [131, 196, 207]. Response-adaptive [131]. retail [167]. Riemannian [33]. river [39, 161]. RNN [123]. RNN-based [123]. Robust [26, 203, 85]. rules [40, 5, 154]. runoff [115].

sample [43, 142]. sampling

[84, 90, 50, 218]. saturated [217]. scales [4]. scheme [161]. schizophrenia [154]. school [196, 220]. schooling [123]. scientific [36]. Score [215, 205]. score-driven [205]. scores [159, 50]. screening [193]. sea [91, 17]. seal [17]. seamless [226]. seasonal [180]. seasonally [146]. Second [150, 63]. second-hand [63]. Second-order [150]. secondary [74]. seed [145]. seemingly [113]. seismic [173]. Selecting [5]. selection [137, 229, 102, 5, 29, 19, 71, 212, 111, 118]. Semi [213, 150, 89, 164, 165]. Semi-parametric [213, 150, 89, 164, 165]. Semiparametric [127, 82]. Senegal [143]. sense [211]. sensed [108]. sensitive [207]. Sensitivity [54, 62, 63, 27]. sensors [184, 98]. separable [191]. sequence [16, 165]. sequences [23]. Sequential [218, 79, 34]. series [104, 146, 83, 195, 209]. service [126]. set [86]. sets [105, 218]. severe [157]. severity [13]. shape [103, 18, 142]. shaped [216]. share [182]. sharing [105, 80]. shock [202]. shooting [65]. short [174]. short-term [174]. shortages [55]. significantly [172]. silencing [45]. Simulating [45]. size [102, 59]. Small [43, 34, 114, 85]. smart [74]. Smirnov [97]. smoke [63]. social [143, 204]. source [172]. sources [165]. soybean [24]. space [108, 169]. space-time [108]. spaced [70]. spaces [142]. Sparse [117, 156, 136]. Spatial [153, 186, 105, 53, 13, 147]. spatially [101, 35]. spatio [168, 152, 160, 199, 191]. spatio-network [152]. spatio-temporal [168, 160, 199, 191]. Spatiotemporal [202, 55, 2]. species [217]. specific [41]. Specification [201]. specificity [62]. spectral [83]. speed [79]. split [176]. spray [44]. spread [168, 24]. squares [206]. Stacked [72]. star [216]. star-shaped [216]. state [168, 160, 169]. statistic [97]. Statistical [49, 206, 53, 201]. statistics [157]. stay [126]. stays [213]. step [222, 218]. stick [147]. stick-breaking [147]. stock [197]. Stopping [195, 40]. storm [179]. strain [181]. strategies [227]. stratification [189]. stratified [41]. streamflow [161]. Strengths [159]. structural [174]. structure [16, 170, 214]. Structured [27, 145]. Student [50]. studies [63]. study [107, 72, 89, 24, 145, 115, 116, 35, 76, 9]. sub [218]. sub-sampling [218]. subdistribution [56]. subgroup [84, 82]. subject [61]. substantiate [36]. subtype [156]. summaries [167]. summarisation [223]. support [70]. surface [103]. surge [179]. survey [8]. surveys [85]. survival [22, 73, 193, 126, 82]. Sustainable [220]. swallow [49]. Swiss [180]. switching [172, 168, 65]. symmetries [132]. synthesis [125]. system [80]. systems [144, 196, 185].

T [110]. T-cell [110]. tailoring [154]. Tanzania [153]. target [171]. targeted [136]. taste [222]. teammate [65]. techniques [86]. technology [201]. teenager [201]. temporal [168, 160, 199, 191]. term [174, 91, 35]. terminal [107, 9]. Terry [153]. test [183, 36, 178, 93, 215, 142]. testing [95, 84, 88, 31, 214]. tests [149, 62, 20, 43]. textural [92]. their [223]. therapies [110]. three [103]. three-dimensional [103]. Threshold [84, 40]. Threshold-based [84, 40]. Time [100, 104, 205, 146, 73, 108, 9, 83, 213, 195, 122, 58, 157, 80, 209]. time-series [195, 209]. time-to-event [213]. time-varying [205, 73, 9, 58, 80]. times [100]. topic [167]. tornado [2]. tracking [144]. trade [114]. trades [114]. trading [11]. trait [113]. trajectories [38]. transaction [97]. transcriptomic [156]. transfer [134]. Transformation [158]. transformations [14]. transformed [191]. transition [21]. trap [176]. treatment [229, 5, 47, 212, 189, 58, 154, 192]. treatment-induced [192]. tree [49, 73, 18]. trees [66, 166]. trend [91]. trendiness [75]. trends [75, 42]. trial [34, 171]. trials [110, 128, 226, 171, 227, 41]. truncated [158]. two [84, 222, 115, 206]. two-field [115]. two-phase [84]. two-step [222]. two-way [206].

UK [74]. Unconventional [197]. unequally [70]. unidimensional [135]. unified [64]. uniform [149]. unifying [187]. unknown [62]. unrelated [113]. unreported [114]. upper [39]. Urban [153]. Urnings [144]. use [114, 201, 50, 42]. Using [20, 143, 86, 38, 106, 130, 44, 5, 112, 120, 135, 108, 17, 221, 207, 18, 185, 10, 57, 98, 157, 80, 190, 146]. Utility [229, 212]. Utility-based [229, 212].

vaccination [143]. vaccine [41]. validity [201]. value [157]. valued [88, 164]. values [40]. variability [65]. Variable [29, 19, 148, 207, 111, 43]. variables [154]. variational [112]. varying [137, 71, 205, 146, 73, 9, 58, 80, 35]. vegetation [108]. vehicles [185]. verbal [96]. via [127, 156, 23, 191]. vine [44]. viruses [180]. visits [48]. visualisation [103, 117]. vivo [181]. volatility [197]. volcanic [7]. volleyball [64]. volume [68, 67, 228, 138, 181]. volumes [114]. Voss [115].

Wald [43]. waste [47]. wastewater [147]. water [47]. way [206]. weight [35]. weighted [130, 72]. weights [50]. well [183, 201, 147]. well-being [201]. while [61]. wind [121, 160, 66, 79]. windstorms [13]. within [104, 98]. wood [195].

York [11].

Zero [168, 20, 116]. zero-inflated [20, 116]. Zero-state [168]. Zoom [4]. Zoom-in [4].

References

Anonymous:2020:IIa

[1] Anonymous. Issue Information. Applied Statistics, 69(1):1–2, January 2020. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gonzalez:2020:ATR

[2] Jonatan A. González, Ute Hahn, and Jorge Mateu. Analysis of tornado reports through replicated spatiotemporal point patterns. Applied Statistics, 69(1): 3–23, January 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Staicu:2020:LDF

[3] Ana-Maria Staicu, Md Nazmul Islam, Raluca Dumitru, and Eric van Heugten. Longitudinal dynamic functional regression. Applied Statistics, 69(1):25–46, January 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pircalabelu:2020:ZJG

[4] Eugen Pircalabelu, Gerda Claeskens, and Lourens J. Waldorp. Zoom-in–out joint graphical lasso for different coarseness scales. Applied Statistics, 69(1):47– 67, January 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dasgupta:2020:SBB

[5] Sayan Dasgupta and Ying Huang. Selecting biomarkers for building optimal treatment selection rules by using kernel machines. Applied Statistics, 69(1):69– 88, January 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Blumke:2020:EPD

[6] Oliver Blümke. Estimating the probability of default for no-default and lowdefault portfolios. Applied Statistics, 69 (1):89–107, January 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wang:2020:BMM

[7] Ting Wang, Matthew Schofield, Mark Bebbington, and Koji Kiyosugi. Bayesian modelling of marked point processes with incomplete records: volcanic eruptions. Applied Statistics, 69(1):109– 130, January 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sahu:2020:FBI

[8] Sujit K. Sahu, Mark R. Bass, Carla Sabariego, Alarcos Cieza, Carolina S. Fellinghauer, and Somnath Chatterji. A full Bayesian implementation of a generalized partial credit model with an application to an international disability survey. Applied Statistics, 69(1):131– 150, January 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2020:TVB

[9] Zheng Li, Vernon M. Chinchilli, and Ming Wang. A time-varying Bayesian joint hierarchical copula model for analysing recurrent events and a terminal event: an application to the Cardiovascular Health Study. Applied Statistics, 69(1):151–166, January 2020. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ruse:2020:IBD

[10] Mareile Große Ruse, Adeline Samson, and Susanne Ditlevsen. Inference for biomedical data by using diffusion models with covariates and mixed effects. Applied Statistics, 69(1):167– 193, January 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Betancourt:2020:MPF

[11] Brenda Betancourt, Abel Rodríguez, and Naomi Boyd. Modelling and prediction of financial trading networks: an application to the New York Mercantile Exchange natural gas futures market. Applied Statistics, 69(1):195– 218, January 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2020:IIb

[12] Anonymous. Issue Information. Applied Statistics, 69(2):221–222, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sharkey:2020:MSE

[13] Paul Sharkey, Jonathan A. Tawn, and Simon J. Brown. Modelling the spatial extent and severity of extreme European windstorms. Applied Statistics, 69(2): 223–250, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atkinson:2020:ATP

[14] Anthony C. Atkinson, Marco Riani, and Aldo Corbellini. The analysis of transformations for profit-and-loss data. Applied Statistics, 69(2):251–275, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chin:2020:EDA

[15] Vincent Chin, David Gunawan, Denzil G. Fiebig, Robert Kohn, and Scott A. Sisson. Efficient data augmentation for multivariate probit models with panel data: an application to general practitioner decision making about contraceptives. Applied Statistics, 69(2):277– 300, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fallaize:2020:BPS

[16] Christopher J. Fallaize, Peter J. Green, Kanti V. Mardia, and Stuart Barber. Bayesian protein sequence and structure alignment. Applied Statistics, 69(2):301– 325, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jullum:2020:ESP

[17] Martin Jullum, Thordis Thorarinsdottir, and Fabian E. Bachl. Estimating seal pup production in the Greenland Sea by using Bayesian hierarchical modelling. Applied Statistics, 69(2):327–352, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liao:2020:EIM

[18] Xiyue Liao and Mary C. Meyer. Estimation and inference in mixed effect regression models using shape constraints, with application to tree height estimation. Applied Statistics, 69(2):353– 375, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Griffin:2020:MED

[19] Jim E. Griffin, Eleni Matechou, Andrew S. Buxton, Dimitrios Bormpoudakis, and Richard A. Griffiths. Modelling environmental DNA data; Bayesian variable selection accounting for false positive and false negative errors. Applied Statistics, 69(2):377– 392, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lipsitz:2020:UCR

[20] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Debajyoti Sinha, Alexander P. Cole, Christian P. Meyer, and Quoc-Dien Trinh. Using Cox regression to develop linear rank tests with zero-inflated clustered data. Applied Statistics, 69(2): 393–411, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeyam:2020:AHT

[21] Anita Jeyam, Rachel McCrea, and Roger Pradel. Assessing heterogeneity in transition propensity in multistate capture–recapture data. Applied Statistics, 69(2):413–427, April 2020. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burke:2020:FPM

[22] Kevin Burke, M. C. Jones, and Angela Noufaily. A flexible parametric modelling framework for survival analysis. Applied Statistics, 69(2):429–457, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Page:2020:BHM

[23] Garritt L. Page, María Xosé Rodríguez-Álvarez, and Dae-Jin Lee. Bayesian hierarchical modelling of growth curve derivatives via sequences of quotient differences. Applied Statistics, 69(2):459– 481, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Marowka:2020:FAB

[24] Maciej Marowka, Gareth W. Peters, Nikolas Kantas, and Guillaume Bagnarosa. Factor-augmented Bayesian cointegration models: a case-study on the soybean crush spread. Applied Statistics, 69(2):483–500, April 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2020:IIc

[25] Anonymous. Issue Information. Applied Statistics, 69(3):501–502, June 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Avery:2020:RAA

[26] Peter Avery, Quentin Clairon, Robin Henderson, C. James Taylor, and Emma Wilson. Robust and adaptive anticoagulant control. Applied Statistics, 69(3): 503–524, June 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhao:2020:SPR

[27] Zhi Zhao and Manuela Zucknick. Structured penalized regression for drug sensitivity prediction. Applied Statistics, 69 (3):525–545, June 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hammon:2020:MIB

[28] Angelina Hammon and Sabine Zinn. Multiple imputation of binary multilevel missing not at random data. Applied Statistics, 69(3):547–564, June 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ghosal:2020:VSF

[29] Rahul Ghosal, Arnab Maity, Timothy Clark, and Stefano B. Longo. Variable selection in functional linear concurrent regression. Applied Statistics, 69 (3):565–587, June 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Klein:2020:JCR

[30] Martin Klein, Tommy Wright, and Jerzy Wieczorek. A joint confidence region for an overall ranking of populations. Applied Statistics, 69(3):589–606, June 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Malinovsky:2020:ODH

[31] Yaakov Malinovsky, Gregory Haber, and Paul S. Albert. An optimal design for hierarchical generalized group testing. Applied Statistics, 69(3):607–621, June 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Denis:2020:NRA

[32] C. Denis, E. Lebarbier, C. Lévy-Leduc, O. Martin, and L. Sansonnet. A novel regularized approach for functional data clustering: an application to milking kinetics in dairy goats. Applied Statistics, 69(3):623–640, June 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Simo:2020:GPL

[33] Amelia Simó, M. Victoria Ibáñez, Irene Epifanio, and Vicent Gimeno. Generalized partially linear models on Riemannian manifolds. Applied Statistics, 69(3): 641–661, June 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chao:2020:BGS

[34] Yan-Cheng Chao, Thomas M. Braun, Roy N. Tamura, and Kelley M. Kidwell. A Bayesian group sequential small n sequential multiple-assignment randomized trial. Applied Statistics, 69(3):663– 680, June 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Warren:2020:SVD

[35] Joshua L. Warren, Thomas J. Luben, and Howard H. Chang. A spatially varying distributed lag model with application to an air pollution and term low birth weight study. Applied Statistics, 69(3):681–696, June 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Held:2020:HMC

[36] Leonhard Held. The harmonic mean χ^{2}- test to substantiate scientific findings. Applied Statistics, 69(3):697–708, June 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2020:IId

[37] Anonymous. Issue information. Applied Statistics, 69(4):709–710, August 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bianchi:2020:LND

[38] Federica Bianchi, Francesco Bartolucci, Stefano Peluso, and Antonietta Mira. Longitudinal networks of dyadic relationships using latent trajectories: evidence from the European interbank market. Applied Statistics, 69(4):711– 739, August 2020. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Mhalla:2020:CME

[39] Linda Mhalla, Valérie Chavez-Demoulin, and Debbie J. Dupuis. Causal mechanism of extreme river discharges in the upper Danube basin network. Applied Statistics, 69(4):741–764, August 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barlow:2020:IEV

[40] Anna Maria Barlow, Chris Sherlock, and Jonathan Tawn. Inference for extreme values under threshold-based stopping rules. Applied Statistics, 69(4):765– 789, August 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sun:2020:HAS

[41] Yanqing Sun, Li Qi, Fei Heng, and Peter B. Gilbert. A hybrid approach for the stratified mark-specific proportional hazards model with missing covariates and missing marks, with application to vaccine efficacy trials. Applied Statistics, 69(4):791–814, August 2020. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stoner:2020:GHE

[42] Oliver Stoner, Gavin Shaddick, Theo Economou, Sophie Gumy, Jessica Lewis, Itzel Lucio, Giulia Ruggeri, and Heather Adair-Rohani. Global household energy model: a multivariate hierarchical approach to estimating trends in the use of polluting and clean fuels for cooking. Applied Statistics, 69(4):815– 839, August 2020. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Ozenne:2020:SSC

[43] Brice Ozenne, Patrick M. Fisher, and Esben Budtz-J⊘rgensen.SmallsamplecorrectionsforW aldtestsinlatentvariablemodels.AppliedStatistics, 69 (4) : 841−−861, August2020.CODENAP ST AG.ISSN0035− 9254(print), 1467 − 9876(electronic).

Coblenz:2020:MFI

[44] Maximilian Coblenz, Simon Holz, Hans-Jörg Bauer, Oliver Grothe, and Rainer Koch. Modelling fuel injector spray characteristics in jet engines by using vine copulas. Applied Statistics, 69(4):863–886, August 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Djordjilovic:2020:SGS

[45] Vera Djordjilović, Monica Chiogna, and Chiara Romualdi. Simulating gene silencing through intervention analysis. Applied Statistics, 69(4):887–907, August 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Grantham:2020:GFG

[46] Neal S. Grantham, Brian J. Reich, Eric B. Laber, Krishna Pacifici, Robert R. Dunn, Noah Fierer, Matthew Gebert, Julia S. Allwood, and Seth A. Faith. Global forensic geolocation with deep neural networks. Applied Statistics, 69(4):909–929, August 2020. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Klanderman:2020:FIC

[47] Molly C. Klanderman, Kathryn B. Newhart, Tzahi Y. Cath, and Amanda S. Hering. Fault isolation for a complex decentralized waste water treatment facility. Applied Statistics, 69(4):931–951, August 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Marra:2020:EBE

[48] Giampiero Marra, Rosalba Radice, and David M. Zimmer. Estimating the binary endogenous effect of insurance on doctor visits by copula-based regression additive models. Applied Statistics, 69(4):953–971, August 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coleman:2020:SIT

[49] Tim Coleman, Lucas Mentch, Daniel Fink, Frank A. La Sorte, David W. Winkler, Giles Hooker, and Wesley M. Hochachka. Statistical inference on tree swallow migrations with random forests. Applied Statistics, 69(4):973– 989, August 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Spagnolo:2020:USW

[50] Francesco Schirripa Spagnolo, Nicola Salvati, Antonella D’Agostino, and Ides Nicaise. The use of sampling weights in M-quantile random-effects regression: an application to Programme for International Student Assessment mathematics scores. Applied Statistics, 69(4):991–1012, August 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2020:IIe

[51] Anonymous. Issue information. Applied Statistics, 69(5):1013–1014, November 2020. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Asar:2020:LME

[52] Özgür Asar, David Bolin, Peter J. Diggle, and Jonas Wallin. Linear mixed effects models for non-Gaussian continuous repeated measurement data. Applied Statistics, 69(5):

1015–1065, November 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Povala:2020:BLI

[53] Jan Povala, Seppo Virtanen, and Mark Girolami. Burglary in London: insights from statistical heterogeneous spatial point processes. Applied Statistics, 69(5):1067–1090, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mathur:2020:SAP

[54] Maya B. Mathur and Tyler J. VanderWeele. Sensitivity analysis for publication bias in meta-analyses. Applied Statistics, 69(5): 1091–1119, November 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ghosal:2020:HME

[55] Soutik Ghosal, Timothy S. Lau, Jeremy Gaskins, and Maiying Kong. A hierarchical mixed effect hurdle model for spatiotemporal count data and its application to identifying factors impacting health professional shortages. Applied Statistics, 69(5):1121–1144, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liu:2020:LPS

[56] Qing Liu, Gong Tang, Joseph P. Costantino, and Chung-Chou H. Chang. Landmark proportional subdistribution hazards models for dynamic prediction of cumulative incidence functions. Applied Statistics, 69(5):1145–1162, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Subedi:2020:CAM

[57] Sanjeena Subedi, Drew Neish, Stephen Bak, and Zeny Feng. Cluster analysis of microbiome data by using mixtures of Dirichletmultinomial regression models. Applied Statistics, 69(5):1163–1187, November 2020. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Spieker:2020:NCC

[58] Andrew J. Spieker, Emily M. Ko, Jason A. Roy, and Nandita Mitra. Nested gcomputation: a causal approach to analysis of censored medical costs in the presence of time-varying treatment. Applied Statistics, 69 (5):1189–1208, November 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McGee:2020:IBE

[59] Glen McGee, Marianthi-Anna Kioumourtzoglou, Marc G. Weisskopf, Sebastien Haneuse, and Brent A. Coull. On the interplay between exposure misclassification and informative cluster size. Applied Statistics, 69 (5):1209–1226, November 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fortunato:2020:OCC

[60] Francesca Fortunato, Laura Anderlucci, and Angela Montanari. One-class classification with application to forensic analysis. Applied Statistics, 69(5):1227–1249, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Karra:2020:AME

[61] Mahesh Karra, David Canning, and Ryoko Sato. Adding measurement error to location data to protect subject confidentiality while allowing for consistent estimation of exposure effects. Applied Statistics, 69(5):1251–1268, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gelman:2020:BAT

[62] Andrew Gelman and Bob Carpenter. Bayesian analysis of tests with unknown specificity and sensitivity. Applied Statistics, 69(5): 1269–1283, November 2020. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2020:CSA

[63] Bo Zhang and Dylan S. Small. A calibrated sensitivity analysis for matched observational studies with application to the effect of secondhand smoke exposure on blood lead levels in children. Applied Statistics, 69(5):1285–1305, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Egidi:2020:BQF

[64] Leonardo Egidi and Ioannis Ntzoufras. A Bayesian quest for finding a unified model for predicting volleyball games. Applied Statistics, 69(5):1307–1336, November 2020. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sandri:2020:MSM

[65] Marco Sandri, Paola Zuccolotto, and Marica Manisera. Markov switching modelling of shooting performance variability and teammate interactions in basketball. Applied Statistics, 69(5):1337–1356, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lang:2020:CRT

[66] Moritz N. Lang, Lisa Schlosser, Torsten Hothorn, Georg J. Mayr, Reto Stauffer, and Achim Zeileis. Circular regression trees and forests with an application to probabilistic wind direction forecasting. Applied Statistics, 69(5):1357–1374, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2020:IAV

[67] Anonymous. Index of authors, volume 69, 2020. Applied Statistics, 69(5):1375–1380, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2020:CV

[68] Anonymous. Contents of volume 69, 2020. Applied Statistics, 69(5):1381–1383, November 2020. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2021:IIa

[69] Anonymous. Issue information. Applied Statistics, 70(1):1–2, January 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Steele:2021:RED

[70] Fiona Steele and Emily Grundy. Random effects dynamic panel models for unequally spaced multivariate categorical repeated measures: an application to child–parent exchanges of support. Applied Statistics, 70(1): 3–23, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heuclin:2021:BVC

[71] Benjamin Heuclin, Frédéric Mortier, Catherine Trottier, and Marie Denis. Bayesian varying coefficient model with selection: an application to functional mapping. Applied Statistics, 70(1):24–50, January 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [137].

Ginestet:2021:SIP

[72] Pablo Gonzalez Ginestet, Ales Kotalik, David M. Vock, Julian Wolfson, and Erin E. Gabriel. Stacked inverse probability of censoring weighted bagging: a case study in the InfCareHIV Register. Applied Statistics, 70(1):51–65, January 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jiang:2021:FES

[73] Shu Jiang, Yijun Xie, and Graham A. Colditz. Functional ensemble survival tree: Dynamic prediction of Alzheimer’s disease progression accommodating multiple time-varying covariates. Applied Statistics, 70(1):66–79, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalair:2021:NPH

[74] Kieran Kalair, Colm Connaughton, and Pierfrancesco Alaimo Di Loro. A non-parametric Hawkes process model of primary and secondary accidents on a UK smart motorway. Applied Statistics, 70(1):80–97, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jensen:2021:QTT

[75] Andreas Kryger Jensen and Claus Thorn Ekstrøm. Quantifying the trendiness of trends. Applied Statistics, 70(1):98–121, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Alfo:2021:MQR

[76] Marco Alfò, Maria Francesca Marino, Maria Giovanna Ranalli, Nicola Salvati, and Nikos Tzavidis. M-quantile regression for multivariate longitudinal data with an application to the Millennium Cohort Study. Applied Statistics, 70(1):122–146, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Santos-Fernandez:2021:CME

[77] Edgar Santos-Fernandez, Erin E. Peterson, Julie Vercelloni, Em Rushworth, and Kerrie Mengersen. Correcting misclassification errors in crowdsourced ecological data: a Bayesian perspective. Applied Statistics, 70(1):147–173, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whitaker:2021:BAD

[78] Gavin A. Whitaker, Ricardo Silva, Daniel Edwards, and Ioannis Kosmidis. A Bayesian approach for determining player abilities in football. Applied Statistics, 70(1):174–201, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zamo:2021:SAP

[79] Michaël Zamo, Liliane Bel, and Olivier Mestre. Sequential aggregation of probabilistic forecasts — application to wind speed ensemble forecasts. Applied Statistics, 70(1):202– 225, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Torti:2021:MTV

[80] Agostino Torti, Alessia Pini, and Simone Vantini. Modelling time-varying mobility flows using function-on-function regression: analysis of a bike sharing system in the city of Milan. Applied Statistics, 70(1):226–247, January 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2021:IIb

[81] Anonymous. Issue information. Applied Statistics, 70(2):249–250, March 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2021:FMS

[82] Lin Zhang and Inyoung Kim. Finite mixtures of semiparametric Bayesian survival kernel machine regressions: application to breast cancer gene pathway subgroup analysis. Applied Statistics, 70(2):251–269, March 2021.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2021:QFA

[83] Ta-Hsin Li. Quantile-frequency analysis and spectral measures for diagnostic checks of time series with nonlinear dynamics. Applied Statistics, 70(2):270–290, March 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2021:TBS

[84] Ying Huang, Juhee Cho, and Youyi Fong. Threshold-based subgroup testing in logistic regression models in two-phase sampling designs. Applied Statistics, 70(2):291–311, March 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:2021:RES

[85] Paul A. Smith, Chiara Bocci, Nikos Tzavidis, Sabine Krieg, and Marc J. E. Smeets. Robust estimation for small domains in business surveys. Applied Statistics, 70(2):312–334, March 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baker:2021:FPB

[86] Evan Baker, Peter Challenor, and Matt Eames. Future proofing a building design using history matching inspired level-set techniques. Applied Statistics, 70(2):335–350, March 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Titman:2021:REM

[87] Andrew C. Titman, Martin J. Wolfsegger, and Thomas F. Jaki. Recurrent events modelling of haemophilia bleeding events. Applied Statistics, 70(2):351–371, March 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lovato:2021:MNH

[88] Ilenia Lovato, Alessia Pini, Aymeric Stamm, Maxime Taquet, and Simone Vantini. Multiscale null hypothesis testing for networkvalued data: analysis of brain networks of patients with autism. Applied Statistics, 70(2):372–397, March 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Josefsson:2021:BSP

[89] Maria Josefsson and Michael J. Daniels. Bayesian semi-parametric G-computation for causal inference in a cohort study with MNAR dropout and death. Applied Statistics, 70 (2):398–414, March 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kim:2021:IBA

[90] Dongah Kim, Krista J. Gile, Honoria Guarino, and Pedro Mateu-Gelabert. Inferring bivariate association from respondent-driven sampling data. Applied Statistics, 70(2):415–433, March 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2021:LTT

[91] Mintaek Lee and Jaechoul Lee. Long-term trend analysis of extreme coastal sea levels with changepoint detection. Applied Statistics, 70(2):434–458, March 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2021:BNM

[92] Xiao Li, Michele Guindani, Chaan S. Ng, and Brian P. Hobbs. A Bayesian nonparametric model for textural pattern heterogeneity. Applied Statistics, 70(2):459–480, March 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stevenson:2021:FYF

[93] Oliver G. Stevenson and Brendon J. Brewer. Finding your feet: a Gaussian process model for estimating the abilities of batsmen in test cricket. Applied Statistics, 70(2):481–506, March 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2021:IIc

[94] Anonymous. Issue information. Applied Statistics, 70(3):509–510, June 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatterjee:2021:MFA

[95] Saptarshi Chatterjee, Shrabanti Chowdhury, and Sanjib Basu. A model-free approach for testing association. Applied Statistics, 70(3): 511–531, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moran:2021:BHF

[96] Kelly R. Moran, Elizabeth L. Turner, David Dunson, and Amy H. Herring. Bayesian hierarchical factor regression models to infer cause of death from verbal autopsy data. Applied Statistics, 70(3):532–557, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhu:2021:CBK

[97] Yingqiu Zhu, Qiong Deng, Danyang Huang, Bingyi Jing, and Bo Zhang. Clustering based on Kolmogorov–Smirnov statistic with application to bank card transaction data. Applied Statistics, 70(3):558–578, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:2021:ADP

[98] Simon A. C. Taylor, Rebecca Killick, Jonathan Burr, and Louise Rogerson. Assessing daily patterns using home activity sensors and within period changepoint detection. Applied Statistics, 70(3):579–595, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Koutra:2021:OBD

[99] Vasiliki Koutra, Steven G. Gilmour, and Ben M. Parker. Optimal block designs for experiments on networks. Applied Statistics, 70(3):596–618, June 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Betz:2021:TMH

[100] Jennifer Betz, Ralf Kellner, and Daniel Rösch. Time matters: How default resolution times impact final loss rates. Applied Statistics, 70(3):619–644, June 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gramatica:2021:BMS

[101] Marco Gramatica, Peter Congdon, and Silvia Liverani. Bayesian modelling for spatially misaligned health areal data: a multiple membership approach. Applied Statistics, 70(3): 645–666, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berusewicz:2021:ESI

[102] Maciej Beresewicz and Dagmara Nikulin. Estimation of the size of informal employment based on administrative records with nonignorable selection mechanism. Applied Statistics, 70(3):667–690, June 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Katina:2021:FDA

[103] Stanislav Katina, Liberty Vittert, and Adrian W. Bowman. Functional data analysis and visualisation of three-dimensional surface shape. Applied Statistics, 70(3):691–713, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gramuglia:2021:CAL

[104] Emanuele Gramuglia, Geir Storvik, and Morten Stakkeland. Clustering and automatic labelling within time series of categorical observations — with an application to marine log messages. Applied Statistics, 70(3):714–732, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lucas:2021:MMS

[105] Tim C. D. Lucas, Anita K. Nandi, Elisabeth G. Chestnutt, Katherine A. Twohig, Suzanne H. Keddie, Emma L. Collins, Rosalind E. Howes, Michele Nguyen, Susan F. Rumisha, Andre Python, Rohan Arambepola, Amelia Bertozzi-Villa, Penelope Hancock, Punam Amratia, Katherine E. Battle, Ewan Cameron, Peter W. Gething, and Daniel J. Weiss. Mapping malaria by sharing spatial information between incidence and prevalence data sets. Applied Statistics, 70(3): 733–749, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cappiello:2021:APD

[106] Lauren Cappiello, Zhiwei Zhang, Changyu Shen, Neel M. Butala, Xinping Cui, and Robert W. Yeh. Adjusting for population differences using machine learning methods. Applied Statistics, 70(3):750–769, June 2021. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ebert:2021:LFP

[107] Anthony Ebert, Ritabrata Dutta, Kerrie Mengersen, Antonietta Mira, Fabrizio Ruggeri, and Paul Wu. Likelihood-free parameter estimation for dynamic queueing networks: Case study of passenger flow in an international airport terminal. Applied Statistics, 70(3):770–792, June 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Johnson:2021:MFR

[108] Margaret C Johnson, Brian J Reich, and Josh M Gray. Multisensor fusion of remotely sensed vegetation indices using space-time dynamic linear models. Applied Statistics, 70(3): 793–812, June 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2021:IId

[109] Anonymous. Issue information. Applied Statistics, 70(4):813–814, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Devlin:2021:PCT

[110] Sean M. Devlin, Alexia Iasonos, and John O’Quigley. Phase I clinical trials in adoptive T-cell therapies. Applied Statistics, 70(4):815– 834, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maity:2021:BCB

[111] Arnab Kumar Maity, Sanjib Basu, and Santu Ghosh. Bayesian criterion-based variable selection. Applied Statistics, 70(4):835–857, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Donnet:2021:ABE

[112] Sophie Donnet and Stéphane Robin. Accelerating Bayesian estimation for network Poisson models using frequentist variational estimates. Applied Statistics, 70(4):858–885, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bottolo:2021:CEB

[113] Leonardo Bottolo, Marco Banterle, Sylvia Richardson, Mika Ala-Korpela, Marjo-Riitta Järvelin, and Alex Lewin. A computationally efficient Bayesian seemingly unrelated regressions model for high-dimensional quantitative trait loci discovery. Applied Statistics, 70(4):886–908, August 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lebacher:2021:CRM

[114] Michael Lebacher, Paul W. Thurner, and Göran Kauermann. Censored regression for modelling small arms trade volumes and its ‘forensic’ use for exploring unreported trades. Applied Statistics, 70(4):909–933, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Roksvaag:2021:TFG

[115] Thea Roksvåg, Ingelin Steinsland, and Kolbjørn Engeland. A two-field geostatistical model combining point and areal observations

— a case study of annual runoff predictions in the Voss area. Applied Statistics, 70(4):934– 960, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shuler:2021:BNA

[116] Kurtis Shuler, Samuel Verbanic, Irene A. Chen, and Juhee Lee. A Bayesian nonparametric analysis for zero-inflated multivariate count data with application to microbiome study. Applied Statistics, 70(4):961–979, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kobak:2021:SRR

[117] Dmitry Kobak, Yves Bernaerts, Marissa A. Weis, Federico Scala, Andreas S. Tolias, and Philipp Berens. Sparse reduced-rank regression for exploratory visualisation of paired multivariate data. Applied Statistics, 70 (4):980–1000, August 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rabier:2021:PHD

[118] Charles-Elie Rabier and Simona Grusea. Prediction in high-dimensional linear models and application to genomic selection under imperfect linkage disequilibrium. Applied Statistics, 70(4):1001–1026, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cui:2021:ARM

[119] Ying Cui, Limin Peng, Yijuan Hu, and HuiChuan J. Lai. Assessing the reproducibility of microbiome measurements based on concordance correlation coefficients. Applied Statistics, 70(4):1027–1048, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Green:2021:IAC

[120] Peter J. Green and Julia Mortera. Inference about complex relationships using peak height data from DNA mixtures. Applied Statistics, 70(4):1049–1082, August 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hardle:2021:PWP

[121] Wolfgang Karl Härdle, Brenda López Cabrera, and Awdesch Melzer. Pricing wind power futures. Applied Statistics, 70(4):1083–1102, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Spence:2021:ACT

[122] Michael A. Spence, Evalyne W. Muiruri, David L. Maxwell, Scott Davis, and Dave Sheahan. The application of continuous-time Markov chain models in the analysis of choice flume experiments. Applied Statistics, 70 (4):1103–1123, August 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Poulos:2021:RBC

[123] Jason Poulos and Shuxi Zeng. RNN-based counterfactual prediction, with an application to homestead policy and public schooling. Applied Statistics, 70(4):1124–1139, August 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2021:IIe

[124] Anonymous. Issue information. Applied Statistics, 70(5):1141–1142, November 2021. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McAlinn:2021:MFB

[125] Kenichiro McAlinn. Mixed-frequency Bayesian predictive synthesis for economic nowcasting. Applied Statistics, 70(5):1143– 1163, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watson:2021:EEH

[126] Samuel I. Watson, Richard J. Lilford, Jianxia Sun, and Julian Bion. Estimating the effect of health service delivery interventions on patient length of stay: a Bayesian survival analysis approach. Applied Statistics, 70(5):1164–1186, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2021:SMA

[127] Jialiang Li, Tonghui Yu, Jing Lv, and Mei-Ling Ting Lee. Semiparametric model averaging prediction for lifetime data via hazards regression. Applied Statistics, 70(5):1187–1209, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Guo:2021:BPI

[128] Beibei Guo, Elizabeth Garrett-Mayer, and Suyu Liu. A Bayesian phase I/II design for cancer clinical trials combining an immunotherapeutic agent with a chemotherapeutic agent. Applied Statistics, 70(5): 1210–1229, November 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Momal:2021:AMA

[129] Raphaëlle Momal, Stéphane Robin, and Christophe Ambroise. Accounting for missing actors in interaction network inference from abundance data. Applied Statistics, 70(5): 1230–1258, November 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chu:2021:DHD

[130] Chenghao Chu and Bingming Yi. Dynamic historical data borrowing using weighted average. Applied Statistics, 70(5):1259–1280, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2021:RAR

[131] Hengtao Zhang and Guosheng Yin. Responseadaptive rerandomization. Applied Statistics, 70(5):1281–1298, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ranciati:2021:FGL

[132] Saverio Ranciati, Alberto Roverato, and Alessandra Luati. Fused graphical lasso for brain networks with symmetries. Applied Statistics, 70(5):1299–1322, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seymour:2021:BNA

[133] Rowland G. Seymour, Theodore Kypraios, Philip D. O’Neill, and Thomas J. Hagenaars. A Bayesian nonparametric analysis of the 2003 outbreak of highly pathogenic avian influenza in the Netherlands. Applied Statistics, 70(5): 1323–1343, November 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Iapteff:2021:RNE

[134] Loïc Iapteff, Julien Jacques, Matthieu Rolland, and Benoit Celse. Reducing the number of experiments required for modelling the hydrocracking process with kriging through Bayesian transfer learning. Applied Statistics, 70(5):1344–1364, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hohberg:2021:BUP

[135] Maike Hohberg, Francesco Donat, Giampiero Marra, and Thomas Kneib. Beyond unidimensional poverty analysis using distributional copula models for mixed ordered-continuous outcomes. Applied Statistics, 70(5):1365–1390, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Song:2021:BSM

[136] Yanyi Song, Xiang Zhou, Jian Kang, Max T. Aung, Min Zhang, Wei Zhao, Belinda L. Needham, Sharon L. R. Kardia, Yongmei Liu, John D. Meeker, Jennifer A. Smith, and Bhramar Mukherjee. Bayesian sparse mediation analysis with targeted penalization of natural indirect effects. Applied Statistics, 70(5): 1391–1412, November 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2021:CBV

[137] Anonymous. Corrigendum: Bayesian varying coefficient model with selection: an application to functional mapping. Applied Statistics, 70(5):1413, November 2021. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [71].

Heuclin:2021:CV

[138] Benjamin Heuclin, Frédéric Mortier, Catherine Trottier, and Marie Denis. Contents of volume 70, 2021. Applied Statistics, 70(1): 1414–1416, November 2021. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2022:IIa

[139] Anonymous. Issue information. Applied Statistics, 71(1):1–2, January 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Karvanen:2022:EML

[140] Juha Karvanen, Mikko Niilo-Rämä, Jan Sarén, and Salme Kärkkäinen. Estimating mean lifetime from partially observed events in nuclear physics. Applied Statistics, 71(1): 3–26, January 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wieditz:2022:CNM

[141] Johannes Wieditz, Yvo Pokern, Dominic Schuhmacher, and Stephan Huckemann. Characteristic and necessary minutiae in fingerprints. Applied Statistics, 71(1):27–50, January 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2022:NST

[142] Ruiyi Zhang, R. Todd Ogden, Martin Picard, and Anuj Srivastava. Nonparametric k-sample test on shape spaces with applications to mitochondrial shape analysis. Applied Statistics, 71(1):51–69, January 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Potter:2022:USC

[143] Gail E. Potter, Nicole Bohme Carnegie, Jonathan D. Sugimoto, Aldiouma Diallo, John C. Victor, Kathleen M. Neuzil, and M. Elizabeth Halloran. Using social contact data to improve the overall effect estimate of a cluster-randomized influenza vaccination program in Senegal. Applied Statistics, 71(1):70– 90, January 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bolsinova:2022:UNM

[144] Maria Bolsinova, Gunter Maris, Abe D. Hofman, Han L. J. van der Maas, and Matthieu J. S. Brinkhuis. Urnings: a new method for tracking dynamically changing parameters in paired comparison systems. Applied Statistics, 71(1):91–118, January 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Perrot-Dockes:2022:ELB

[145] M. Perrot-Dockès, C. Lévy-Leduc, and L. Rajjou. Estimation of large block structured covariance matrices: Application to ‘multi-omic’ approaches to study seed quality. Applied Statistics, 71(1):119–147, January 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2022:DMM

[146] Guowen Huang, Patrick E. Brown, Sze Hang Fu, and Hwashin Hyun Shin. Daily mortality/morbidity and air quality: Using multivariate time series with seasonally varying covariances. Applied Statistics, 71(1):148–174, January 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Warren:2022:DKS

[147] Joshua L. Warren, Jiachen Cai, Nicholaus P. Johnson, and Nicole C. Deziel. A discrete kernel stick-breaking model for detecting spatial boundaries in hydraulic fracturing wastewater disposal well placement across Ohio. Applied Statistics, 71(1):175–193, January 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holmes:2022:PGD

[148] John B. Holmes, Matthew R. Schofield, and Richard J. Barker. Pólya-gamma data augmentation and latent variable models for multivariate binomial data. Applied Statistics, 71(1):194–218, January 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cousido-Rocha:2022:MCP

[149] Marta Cousido-Rocha, Jacobo de Uña-Álvarez, and Sebastian Döhler. Multiple comparison procedures for discrete uniform and homogeneous tests. Applied Statistics, 71 (1):219–243, January 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hessellund:2022:SOS

[150] Kristian Bjørn Hessellund, Ganggang Xu, Yongtao Guan, and Rasmus Waagepetersen. Second-order semi-parametric inference for multivariate log Gaussian Cox processes. Applied Statistics, 71(1):244–268, January 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2022:IIb

[151] Anonymous. Issue information. Applied Statistics, 71(2):269–270, March 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gerogiannis:2022:BSN

[152] George Gerogiannis, Mark Tranmer, Duncan Lee, and Thomas Valente. A Bayesian spationetwork model for multiple adolescent adverse health behaviours. Applied Statistics, 71(2):271–287, March 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seymour:2022:BSB

[153] Rowland G. Seymour, David Sirl, Simon P. Preston, Ian L. Dryden, Madeleine J. A. Ellis, Bertrand Perrat, and James Goulding. The Bayesian Spatial Bradley–Terry model: Urban deprivation modelling in Tanzania. Applied Statistics, 71(2):288–308, March 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wu:2022:RTV

[154] Jiacheng Wu, Nina Galanter, Susan M. Shortreed, and Erica E. M. Moodie. Ranking tailoring variables for constructing individualized treatment rules: an application to schizophrenia. Applied Statistics, 71(2):309–330, March 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pourmohamad:2022:GPM

[155] Tony Pourmohamad, Cristian M. Oliva Avilés, and Robert Richardson. Gaussian process modeling for dissolution curve comparisons. Applied Statistics, 71(2):331–351, March 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Meng:2022:OGS

[156] Lingsong Meng, Dorina Avram, George Tseng, and Zhiguang Huo. Outcome-guided sparse K-means for disease subtype discovery via integrating phenotypic data with highdimensional transcriptomic data. Applied Statistics, 71(2):352–375, March 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomas:2022:RTP

[157] Maud Thomas and Holger Rootzén. Realtime prediction of severe influenza epidemics using extreme value statistics. Applied Statistics, 71(2):376–394, March 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Qian:2022:TMB

[158] Jing Qian, Sy Han Chiou, and Rebecca A. Betensky. Transformation model based regression with dependently truncated and independently censored data. Applied Statistics, 71(2):395–416, March 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Merlo:2022:QMH

[159] Luca Merlo, Lea Petrella, and Nikos Tzavidis. Quantile mixed hidden Markov models for multivariate longitudinal data: an application to children’s Strengths and Difficulties Questionnaire scores. Applied Statistics, 71(2):417– 448, March 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2022:FHF

[160] Huang Huang, Stefano Castruccio, and Marc G. Genton. Forecasting high-frequency spatio-temporal wind power with dimensionally reduced echo state networks. Applied Statistics, 71(2):449–466, March 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Park:2022:LSS

[161] Seoncheol Park and Hee-Seok Oh. Lifting scheme for streamflow data in river networks. Applied Statistics, 71(2):467–490, March 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liu:2022:E

[162] Suyu Liu and Ying Yuan. Erratum. Applied Statistics, 71(2):491–492, March 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2022:IIc

[163] Anonymous. Issue information. Applied Statistics, 71(3):493–494, June 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rao:2022:SPI

[164] Yao Rao, David Harris, and Brendan McCabe. A semi-parametric integer-valued autoregressive model with covariates. Applied Statistics, 71(3):495–516, June 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Xi:2022:ISH

[165] Xiaoyue Xi, Simon E. F. Spencer, Matthew Hall, M. Kate Grabowski, Joseph Kagaayi, Oliver Ratmann, on behalf of Rakai Health Sciences Program, and Pangea-Hiv. Inferring the sources of HIV infection in Africa from deep-sequence data with semi-parametric Bayesian Poisson flow models. Applied Statistics, 71(3):517–540, June 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wang:2022:RRC

[166] Bingkai Wang, Brian S. Caffo, Xi Luo, Chin-Fu Liu, Andreia V. Faria, Michael I. Miller, Yi Zhao, and for the Alzheimer’s Disease Neuroimaging Initiative*. Regularized regression on compositional trees with application to MRI analysis. Applied Statistics, 71(3):541– 561, June 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carrasco:2022:PSG

[167] Mariflor Vega Carrasco, Ioanna Manolopoulou, Jason O’Sullivan, Rosie Prior, and Mirco Musolesi. Posterior summaries of grocery retail topic models: Evaluation, interpretability and credibility. Applied Statistics, 71(3):562–588, June 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Douwes-Schultz:2022:ZSC

[168] Dirk Douwes-Schultz and Alexandra M. Schmidt. Zero-state coupled Markov switching count models for spatio-temporal infectious disease spread. Applied Statistics, 71(3): 589–612, June 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kreuzer:2022:BNL

[169] Alexander Kreuzer, Luciana Dalla Valle, and Claudia Czado. A Bayesian non-linear state space copula model for air pollution in Beijing. Applied Statistics, 71(3):613–638, June 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Guha:2022:PPB

[170] Subharup Guha, Rex Jung, and David Dunson. Predicting phenotypes from brain connection structure. Applied Statistics, 71(3): 639–668, June 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2022:GTE

[171] Fan Li, Ashley L. Buchanan, and Stephen R. Cole. Generalizing trial evidence to target populations in non-nested designs: Applications to AIDS clinical trials. Applied Statistics, 71(3):669–697, June 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berentsen:2022:MCC

[172] Geir D. Berentsen, Jan Bulla, Antonello Maruotti, and Bård Støve. Modelling clusters of corporate defaults: Regime-switching models significantly reduce the contagion source. Applied Statistics, 71(3):698–722, June 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jupp:2022:ASO

[173] P. E. Jupp, I. B. J. Goudie, R. A. Batchelor, and R. J. B. Goudie. Assessing seismic origin of geological features by fitting equidistant parallel lines. Applied Statistics, 71(3): 723–738, June 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gamerman:2022:DSE

[174] Dani Gamerman, Luigi Ippoliti, and Pasquale Valentini. A dynamic structural equation approach to estimate the short-term effects of air pollution on human health. Applied Statistics, 71(3):739–769, June 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2022:IId

[175] Anonymous. Issue information. Applied Statistics, 71(4):771–772, August 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Koshy:2022:EBA

[176] Aranya Koshy and Shahin Tavakoli. Exploring British accents: Modelling the trap–bath split with functional data analysis. Applied Statistics, 71(4):773–805, August 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Petrovich:2022:HIF

[177] Justin Petrovich, Matthew Reimherr, and Carrie Daymont. Highly irregular functional generalized linear regression with electronic health records. Applied Statistics, 71(4):806– 833, August 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jersakova:2022:BIC

[178] Radka Jersakova, James Lomax, James Hetherington, Brieuc Lehmann, George Nicholson, Mark Briers, and Chris Holmes. Bayesian imputation of COVID-19 positive test counts for nowcasting under reporting lag. Applied Statistics, 71(4):834–860, August 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ma:2022:MCM

[179] Pulong Ma, Georgios Karagiannis, Bledar A. Konomi, Taylor G. Asher, Gabriel R. Toro, and Andrew T. Cox. Multifidelity computer model emulation with high-dimensional output: an application to storm surge. Applied Statistics, 71(4):861–883, August 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ranjbar:2022:MES

[180] Setareh Ranjbar, Eva Cantoni, Valérie Chavez-Demoulin, Giampiero Marra, Rosalba Radice, and Katia Jaton. Modelling the extremes of seasonal viruses and hospital congestion: the example of flu in a Swiss hospital. Applied Statistics, 71(4):884–905, August 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lazarus:2022:ICM

[181] Alan Lazarus, Hao Gao, Xiaoyu Luo, and Dirk Husmeier. Improving cardio-mechanic inference by combining in vivo strain data with ex vivo volume–pressure data. Applied Statistics, 71(4):906–931, August 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mastrantonio:2022:MMM

[182] Gianluca Mastrantonio. The modelling of movement of multiple animals that share behavioural features. Applied Statistics, 71 (4):932–950, August 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Botsas:2022:BMR

[183] Themistoklis Botsas, Jonathan A. Cumming, and Ian H. Jermyn. A Bayesian multi-region radial composite reservoir model for deconvolution in well test analysis. Applied Statistics, 71(4):951–968, August 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jacques:2022:ACS

[184] Julien Jacques and Sanja Samardžić. Analysing cycling sensors data through ordinal logistic regression with functional covariates. Applied Statistics, 71(4):969–986, August 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Min:2022:RAA

[185] Jie Min, Yili Hong, Caleb B. King, and William Q. Meeker. Reliability analysis of artificial intelligence systems using recurrent events data from autonomous vehicles. Applied Statistics, 71(4):987–1013, August 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:2022:NNG

[186] Tilman M. Davies, Sudipto Banerjee, Adam P. Martin, and Rose E. Turnbull. A nearestneighbour Gaussian process spatial factor model for censored, multi-depth geochemical data. Applied Statistics, 71(4):1014–1043, August 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Eletti:2022:UFF

[187] Alessia Eletti, Giampiero Marra, Manuela Quaresma, Rosalba Radice, and Francisco Javier Rubio. A unifying framework for flexible excess hazard modelling with applications in cancer epidemiology. Applied Statistics, 71(4):1044–1062, August 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2022:IIe

[188] Anonymous. Issue information. Applied Statistics, 71(5):1063–1064, November 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schnell:2022:PSA

[189] Patrick M. Schnell, Richard Baumgartner, Shahrul Mt-Isa, and Vladimir Svetnik. A principal stratification approach to estimating the effect of continuing treatment after observing early outcomes. Applied Statistics, 71 (5):1065–1084, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yang:2022:CMB

[190] Yalei Yang, Hao Gao, Colin Berry, David Carrick, Aleksandra Radjenovic, and Dirk Husmeier. Classification of myocardial blood flow based on dynamic contrast-enhanced magnetic resonance imaging using hierarchical Bayesian models. Applied Statistics, 71(5):1085–1115, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prates:2022:NSS

[191] Marcos O. Prates, Douglas R. M. Azevedo, Ying C. MacNab, and Michael R. Willig. Nonseparable spatio-temporal models via transformed multivariate Gaussian Markov random fields. Applied Statistics, 71(5):1116–1136, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2022:APD

[192] Cuihong Zhang, Jing Ning, Steven H. Belle, Robert H. Squires, Jianwen Cai, and Ruosha Li. Assessing predictive discrimination performance of biomarkers in the presence of treatment-induced dependent censoring. Applied Statistics, 71(5):1137–1157, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Qiu:2022:DDS

[193] Peihua Qiu and Lu You. Dynamic disease screening by joint modelling of survival and longitudinal data. Applied Statistics, 71(5): 1158–1180, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Molinari:2022:BNM

[194] Marco Molinari, Andrea Cremaschi, Maria De Iorio, Nishi Chaturvedi, Alun D. Hughes, and Therese Tillin. Bayesian nonparametric modelling of multiple graphs with an application to ethnic metabolic differences. Applied Statistics, 71(5):1181–1204, November 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shang:2022:STD

[195] Han Lin Shang, Jiguo Cao, and Peijun Sang. Stopping time detection of wood panel compression: a functional time-series approach. Applied Statistics, 71(5):1205–1224, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jin:2022:MNI

[196] Ick Hoon Jin, Minjeong Jeon, Michael Schweinberger, Jonghyun Yun, and Lizhen Lin. Multilevel network item response modelling for discovering differences between innovation and regular school systems in Korea. Applied Statistics, 71(5):1225–1244, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lacava:2022:UPE

[197] Demetrio Lacava, Giampiero M. Gallo, and Edoardo Otranto. Unconventional policies effects on stock market volatility: The MAP approach. Applied Statistics, 71(5):1245–1265, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seaman:2022:NCD

[198] Shaun R. Seaman, Pantelis Samartsidis, Meaghan Kall, and Daniela De Angelis. Nowcasting COVID-19 deaths in England by age and region. Applied Statistics, 71(5): 1266–1281, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pasanen:2022:BST

[199] Tiia-Maria Pasanen, Miikka Voutilainen, Jouni Helske, and Harri Högmander. A Bayesian spatio-temporal analysis of markets during the Finnish 1860s famine. Applied Statistics, 71(5):1282–1302, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2022:HGM

[200] Jiaqi Zhang, Xinyan Fan, Yang Li, and Shuangge Ma. Heterogeneous graphical model for non-negative and non-Gaussian PM2.5 data. Applied Statistics, 71(5):1303–1329, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Semken:2022:SAT

[201] Christoph Semken and David Rossell. Specification analysis for technology use and teenager well-being: Statistical validity and a Bayesian proposal. Applied Statistics, 71(5):1330–1355, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stindl:2022:SEM

[202] Tom Stindl and Feng Chen. Spatiotemporal ETAS model with a renewal main-shock arrival process. Applied Statistics, 71(5): 1356–1380, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Riani:2022:RCA

[203] Marco Riani, Anthony C. Atkinson, Francesca Torti, and Aldo Corbellini. Robust correspondence analysis. Applied Statistics, 71(5): 1381–1401, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:2022:NHP

[204] Owen G. Ward, Jing Wu, Tian Zheng, Anna L. Smith, and James P. Curley. Network Hawkes process models for exploring latent hierarchy in social animal interactions. Applied Statistics, 71(5):1402–1426, November 2022. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holy:2022:MTV

[205] Vladimír Holý and Jan Zouhar. Modelling time-varying rankings with autoregressive and score-driven dynamics. Applied Statistics, 71 (5):1427–1450, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

elBouhaddani:2022:SIH

[206] Said el Bouhaddani, Hae-Won Uh, Geurt Jongbloed, and Jeanine Houwing-Duistermaat. Statistical integration of heterogeneous omics data: Probabilistic two-way partial least squares (PO2PLS). Applied Statistics, 71(5): 1451–1470, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2022:IAS

[207] Shen-Ming Lee, Truong-Nhat Le, Phuoc-Loc Tran, and Chin-Shang Li. Investigating the association of a sensitive attribute with a random variable using the Christofides generalised randomised response design and Bayesian methods. Applied Statistics, 71(5): 1471–1502, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Erciulescu:2022:MBA

[208] Andreea L. Erciulescu and Jean D. Opsomer. A model-based approach to predict employee compensation components. Applied Statistics, 71(5):1503–1520, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yang:2022:NPB

[209] Daewon Yang, Taeryon Choi, Eric Lavigne, and Yeonseung Chung. Non-parametric Bayesian covariate-dependent multivariate functional clustering: an application to timeseries data for multiple air pollutants. Applied Statistics, 71(5):1521–1542, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mancini:2022:EEC

[210] Tullio Mancini, Hector Calvo-Pardo, and Jose Olmo. Environmental Engel curves: a neural network approach. Applied Statistics, 71 (5):1543–1568, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zafar:2022:MDS

[211] Schyan Zafar and Geoff K. Nicholls. Measuring diachronic sense change: New models and Monte Carlo methods for Bayesian inference. Applied Statistics, 71(5):1569–1604, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2022:UBB

[212] Juhee Lee, Peter F. Thall, Bora Lim, and Pavlos Msaouel. Utility-based Bayesian personalized treatment selection for advanced breast cancer. Applied Statistics, 71(5):1605–1622, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2022:SPT

[213] Yang Li, Hao Liu, Xiaoshen Wang, and Wanzhu Tu. Semi-parametric time-to-event modelling of lengths of hospital stays. Applied Statistics, 71(5):1623–1647, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sewell:2022:LNS

[214] Daniel K. Sewell. Leveraging network structure to improve pooled testing efficiency. Applied Statistics, 71(5):1648–1662, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Xu:2022:STA

[215] Huiping Xu, Xiaochun Li, Zuoyi Zhang, and Shaun Grannis. Score test for assessing the conditional dependence in latent class models and its application to record linkage. Applied Statistics, 71(5):1663–1687, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Director:2022:CMP

[216] Hannah M. Director and Adrian E. Raftery. Contour models for physical boundaries enclosing star-shaped and approximately starshaped polygons. Applied Statistics, 71(5): 1688–1720, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Flint:2022:SPI

[217] Ian Flint, Nick Golding, Peter Vesk, Yan Wang, and Aihua Xia. The saturated pairwise interaction Gibbs point process as a joint species distribution model. Applied Statistics, 71(5):1721–1752, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wang:2022:SOS

[218] Feifei Wang, Danyang Huang, Tianchen Gao, Shuyuan Wu, and Hansheng Wang. Sequential one-step estimator by sub-sampling for customer churn analysis with massive data sets. Applied Statistics, 71(5):1753–1786, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cortese:2022:EEM

[219] Giuliana Cortese and Thomas H. Scheike. Efficient estimation of the marginal mean of recurrent events. Applied Statistics, 71(5):

1787–1821, November 2022. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dharamshi:2022:BME

[220] Ameer Dharamshi, Bilal Barakat, Leontine Alkema, and Manos Antoninis. A Bayesian model for estimating Sustainable Development Goal indicator 4.1.2: School completion rates. Applied Statistics, 71(5):1822–1864, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krennmair:2022:FDP

[221] Patrick Krennmair and Timo Schmid. Flexible domain prediction using mixed effects random forests. Applied Statistics, 71(5):1865–1894, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nikiforova:2022:OAC

[222] Nedka Dechkova Nikiforova, Rossella Berni, and Jesús Fernando López-Fidalgo. Optimal approximate choice designs for a two-step coffee choice, taste and choice again experiment. Applied Statistics, 71(5):1895–1917, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heaton:2022:NPC

[223] Timothy J. Heaton. Non-parametric calibration of multiple related radiocarbon determinations and their calendar age summarisation. Applied Statistics, 71(5):1918–1956, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stephenson:2022:DMD

[224] Briana J. K. Stephenson, Amy H. Herring, and Andrew F. Olshan. Derivation of maternal dietary patterns accounting for regional heterogeneity. Applied Statistics, 71(5):1957–1977, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2022:BML

[225] Hanwen Huang. Bayesian multi-level mixedeffects model for influenza dynamics. Applied Statistics, 71(5):1978–1995, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jimenez:2022:CCA

[226] José L. Jiménez and Mourad Tighiouart. Combining cytotoxic agents with continuous dose levels in seamless phase I-II clinical trials. Applied Statistics, 71(5):1996–2013, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ouma:2022:BMS

[227] Luke O. Ouma, Michael J. Grayling, James M. S. Wason, and Haiyan Zheng. Bayesian modelling strategies for borrowing of information in randomised basket trials. Applied Statistics, 71(5):2014–2037, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2022:CV

[228] Anonymous. Contents of volume 71, 2022. Applied Statistics, 71(5):2038–2041, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2022:UBB

[229] Anonymous. Utility-based Bayesian personalized treatment selection for advanced breast cancer. Applied Statistics, 71(5):O1, November 2022. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).